OpenStudy (anonymous):

find the value of k so that the given points are 6/5 untis apart (k,6), (-2,-6)

OpenStudy (anonymous):

Do you know the distance formula between two points?

OpenStudy (anonymous):

\[\sqrt{(B-b)^2-(A-a)^2}\] for two points (B,A) (b,a)

OpenStudy (anonymous):

just sub in the points and let the distance equal 6/5 then solve.

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

\[\sqrt{(k+(-2))^2-(6+(-6))^2}=6/5\] can you solve the above?